A movie theater had a certain number of tickets to give away. Five people got the tickets. The first got one-third of the tickets, the second got one-fourth of the tickets, and the third got one-fifth of the tickets. The fourth person got eight tickets, and there were five tickets left for the fifth person. Find the total number of tickets given away.


Found 2 solutions by richwmiller, mananth:
Answer by richwmiller(17219)


5 people
1/3x went to the first
1/4x went to the second
1/5x went to the third
8 went to the fourth
5 for the fifth
This will work out okay
x*(1/3+1/4+1/5)+13=x
x=60
check
1/3*60=20
1/4*60=15
1/5*60=12
13
total 60 people

Answer by mananth(16949)

A movie theater had a certain number of tickets to give away. Five people got the tickets. The first got one-third of the tickets, the second got one-fourth of the tickets, and the third got one-fifth of the tickets. The fourth person got eight tickets, and there were five tickets left for the fifth person. Find the total number of tickets given away.

Let the number of tickets available be x
1/3 x +1/4 x +1/5 x +8+5 =x
20x+15x+12x / 60 +13 =x
47x/60 -x+13=0
47x-60x/60 =-13
-13x=-780
x=60 Number of tickets given away